### v1

# Stabilizing black-box model selection with the inflated argmax

Abstract
Model selection is the process of choosing from a class of candidate models given data. For instance, methods such as the LASSO and sparse identification of nonlinear dynamics (SINDy) formulate model selection as finding a sparse solution to a linear system of equations determined by training data. However, absent strong assumptions, such methods are highly unstable: if a single data point is removed from the training set, a different model may be selected. This paper presents a new approach to stabilizing model selection that leverages a combination of bagging and an “inflated” argmax operation. Our method selects a small collection of models that all fit the data, and it is stable in that, with high probability, the removal of any training point will result in a collection of selected models that overlaps with the original collection. In addition to developing theoretical guarantees, we illustrate this method in (a) a simulation in which strongly correlated covariates make standard LASSO model selection highly unstable and (b) a Lotka–Volterra model selection problem focused on identifying how competition in an ecosystem influences species’ abundances. In both settings, the proposed method yields stable and compact collections of selected models, outperforming a variety of benchmarks.
1 Introduction
Scientists often choose between many models to identify key aspects of a given phenomenon. In areas such as economics and biology, models are typically idealized (there is little sense in talking about a ‘true’ model) but also interpretable. The final selection is pragmatically treated as getting at some important truth of the structure of the underlying phenomenon.
A necessary condition for trustworthy model selection procedures is stability, i.e., the selection should not be too sensitive to small changes in the data (Yu,, 2013). With limited information, rich model classes, or complex phenomena, principled methods may not be able to reliably and definitively extract a single best model.
Bagging (Breiman, 1996a, ; Breiman, 1996b, ) has been widely applied to stabilize model selection procedures, see § 3 for examples. When the base procedure is unstable, bagging places near-equal weight on models with similar fits to the data. Bagging can thus express model uncertainty, but, as we show empirically in § 5, selecting the most frequent model among multiple close contenders can still be quite unstable.
1.1 Our contributions
In this work, we allow model selection procedures to return a set of candidate models, conveying uncertainty about which model is best while returning a single model as often as possible. We formulate a notion of model selection stability based on the idea that our claims to which model is best should remain logically consistent when we drop a small amount of data.
We develop a framework for stabilizing model selection procedures by combining bagged model selection with the inflated argmax (Soloff et al., 2024a, ). We give a theoretical guarantee for model selection stability associated with our framework. We show empirically that the inflated argmax can improve model stability, particularly in the presence of highly correlated covariates, a setting known to yield instabilities in standard model selection algorithms.
2 Our framework
A model selection procedure is a map $\mathcal{M}$ from data $\mathcal{D}$ to a set of selected models. The output $\hat{M}=\mathcal{M}(\mathcal{D})$ is a nonempty subset of all candidate models, denoted $M^{+}$. If $\hat{M}$ is always a singleton, $\mathcal{M}$ is called simple.
Example. In variable selection (a special case of model selection), the data consists of $n$ variable-response pairs $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$, where $x_{i}\in\mathbb{R}^{d}$. We seek to identify relevant variables, so the model class $M^{+}=2^{[d]}$ is the power set of $[d]=\{1,\ldots,d\}$.
The model class need not consist of subsets of variables. Each $m\in M^{+}$ could be a different learning rule $f:\mathcal{X}\to\mathcal{Y}$, or a probability distribution $P$ meant to describe the data $\mathcal{D}$. For instance, in graphical model selection, $M^{+}$ is the set of all graphs $G$ on $d$ vertices, where an edge $(j,k)$ in $G$ encodes conditional dependence between variables $j$ and $k$ (Drton and Perlman,, 2004; Friedman et al.,, 2008). In dynamical system identification, $M^{+}$ is a set of possible differential equations governing the dynamics underlying noisy data (Brunton et al.,, 2016). In econometrics, $M^{+}$ could be a collection of competing time series forecasts (Masini et al.,, 2023). To keep our discussion general, the model class is an abstract, countable set $M^{+}$.
We decompose the procedure $\mathcal{M}$ into two stages
$\mathcal{M}=\mathcal{S}\circ\mathcal{A}$ (1)
where $\mathcal{A}$ maps data $\mathcal{D}$ to weights $\hat{w}=\mathcal{A}(\mathcal{D})$ on the model class $M^{+}$, i.e., $\hat{w}_{m}$ is the weight assigned to model $m\in M^{+}$ and is higher for models that ${\cal A}$ deems better fit the data. For instance, in Bayesian model selection, $\hat{w}_{m}$ corresponds to the posterior probability of $m$ given the data $\mathcal{D}$. The selection rule $\mathcal{S}$ takes these weights $\hat{w}$ as input and returns a non-empty set of selected models $\mathcal{S}(\hat{w})\subseteq M^{+}$.
Even if $\mathcal{M}$ is only accessible as a black-box, we can define $\mathcal{S}$ and $\mathcal{A}$ such that (1) holds. Specifically, if
$(\mathcal{A}_{\mathcal{M}}(\mathcal{D}))_{m}:=\frac{1}{|\mathcal{M}(\mathcal{D})|}\mathbbm{1}\{m\in\mathcal{M}(\mathcal{D})\},$ (2)
then we can interpret $\mathcal{M}$ as $\mathop{\!\arg\!\max}\circ\,\mathcal{A}_{\mathcal{M}}$ – i.e., here ${\cal S}=\mathop{\!\arg\!\max}$ is selecting models returned by ${\cal A}$ with the largest weight.
We now provide a brief overview of our pipeline for stabilizing a model weighting algorithm $\mathcal{A}$.
  1. 1.
    First, we bag the weighting algorithm $\mathcal{A}$, i.e., average the weights fit on different random samples from $\mathcal{D}$. We fully define bagging in Algorithm 1.
  2. 2.
    Next, we pass through the bagged weights $\hat{w}$ through the inflated argmax, which selects a set of near-maximizers of $\hat{w}$ that is robust to small perturbations of the weights. We define the inflated argmax in Algorithm 2.
In § 2.3, we provide a stability guarantee for this pipeline that can be applied to any algorithm $\mathcal{A}$ and any data set $\mathcal{D}$. We first discuss our pipeline in detail.
2.1 Bagged model selection
The output of the weighting algorithm $\mathcal{A}$ belongs to the probability simplex $\Delta_{|M^{+}|-1}$. Bagging averages an ensemble of weight vectors fit on different subsets of the data — see Algorithm 1. We denote the resulting bagged algorithm by $\widetilde{\mathcal{A}}_{K,B}$, where $K$ is the number of samples in a bag and $B$ is the number of bags. Sampling with replacement is known as bagging, and sampling without replacement is known as subbagging.
For simple model selection procedures, the ensemble weight $\hat{w}_{m}$ simply counts the fraction of bags wherein model $m$ was selected. Even if the model class $M^{+}$ is countably infinite, the weights $\hat{w}$ will have at most $B$ nonzero entries, so bagging is still tractable.
Algorithm 1 Bagged model selection
0:  Model weighting algorithm $\mathcal{A}$; data $\mathcal{D}$; ensemble size $B$; bag size $K$
  for $b=1,\ldots,B$ do
     Construct $\mathcal{D}^{b}$ by sampling uniformly (with or without replacement) $K$ times from $\mathcal{D}$
     Compute model weights $\hat{w}^{b}=\mathcal{A}(\mathcal{D}^{b})$
  end for
  Ensemble weights $\hat{w}=\frac{1}{B}\sum_{b=1}^{B}\hat{w}^{b}$
Algorithm 2 Inflated Argmax
0:  Weights $\hat{w}$; inflation tolerance $\varepsilon$
  Solve for $c$ such that
$\left(\sum_{m\in M^{+}}(\hat{w}_{m}-c)_{+}\right)^{2}+\sum_{m\in M^{+}}(\hat{w}_{m}-c)^{2}_{+}=\varepsilon^{2}$
  $\left\{m\in M^{+}:\hat{w}_{m}>c+\frac{\varepsilon}{\sqrt{2}}-\sum_{m^{\prime}}(\hat{w}_{m^{\prime}}-c)_{+}\right\}$
In Bayesian inference, $\mathcal{A}(\mathcal{D})$ is a posterior distribution over models $m\in M^{+}$, so $\widetilde{\mathcal{A}}_{K,B}(\mathcal{D})$ is the average of $B$ posteriors fit to subsets of size $K$. This method, known as BayesBag, has been previously applied to Bayesian model selection (Huggins and Miller,, 2023).
Soloff et al., 2024b show that bagged weights are stable, meaning that dropping a single sample cannot change the bagged weights $\hat{w}$ by much in the Euclidean norm. However, in order to translate this into any kind of guarantee on the selected model, a new selection rule $\mathcal{S}$ is needed, as we discuss in the next section.
2.2 The inflated argmax
Soloff et al., 2024a introduced the inflated argmax to guarantee stability in the context of multiclass classification. The inflated argmax, denoted by argmaxε, is a selection rule that will allow us to decide which subset $M$ of $M^{+}$ to return.
Definition 2.1 (Inflated argmax).
For $m\in M^{+}$, let
$R_{m}^{\varepsilon}=\left\{w\in\Delta_{|M^{+}|-1}:w_{m}\geq\max_{m^{\prime}\neq m}w_{m^{\prime}}+\frac{\varepsilon}{\sqrt{2}}\right\}.$ (3)
For any $w\in\Delta_{|M^{+}|-1}$ and $\varepsilon>0$, the inflated argmax is defined as
$\textnormal{argmax}^{\varepsilon}(w):=\left\{m\in M^{+}:\textnormal{dist}(w,R_{m}^{\varepsilon})<\varepsilon\right\},$ (4)
where dist$(w,R_{m}^{\varepsilon})=\inf_{q\in R_{m}^{\varepsilon}}||q-w||.$
An intuitive explanation of Equation 4 is that $m\in\text{argmax}^{\varepsilon}(w)$ indicates that $m$ would have the largest weight in $w$ by some margin if the weights $w$ were slightly perturbed, where both the size of the margin and the weight perturbations scale with $\varepsilon$. This definition naturally allows for the number of returned models to adapt to the uncertainty in the model selection: for example, if the top two competing models are close in probability, then the inflated argmax will return both models for some choice of small $\varepsilon$. On the other hand, if the top two competing models have a large enough gap in their two weights, the inflated argmax will only return the top model for a choice of small $\varepsilon$. We present a visual representation of the inflated argmax in the supplement.
2.3 Stability guarantee
In this section, we define stability in the context of model selection and state a theorem guaranteeing the stability of our method.
To assess the stability of $\mathcal{M}$, we measure how the output of $\mathcal{M}$ changes with small perturbations applied to $\mathcal{D}$. We adapt the definition of stability in Soloff et al., 2024a to the context of model selection.
Definition 2.2.
A procedure $\mathcal{M}$ has model selection stability $\delta$ at sample size $n$ if, for all datasets $\mathcal{D}$ with $n$ samples,
$\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\{\hat{M}\cap\hat{M}^{\setminus i}=\varnothing\}\leq\delta,$ (5)
where $\hat{M}=\mathcal{M}(\mathcal{D})$ and $\hat{M}^{\setminus i}=\mathcal{M}(\mathcal{D}^{\setminus i})$ for each $i\in[n]$.
In Definition 2.2, $\mathcal{D}^{\setminus i}$ refers to dataset $\mathcal{D}$ with the $i$-th sample removed. In plain language, the inequality in (5) quantifies stability in terms of whether the selected models overlap when we drop a single observation at random.
Theorem 1.
For any model selection procedure $\mathcal{S}\circ\mathcal{A}$, our method $\textnormal{argmax}^{\varepsilon}\circ\widetilde{A}_{K,B}$ satisfies model selection stability at instability level
$\delta=\frac{1}{\varepsilon^{2}}\left(1-\frac{1}{|M^{+}|}\right)\left(\frac{1}{n-1}\frac{\rho}{1-\rho}+\frac{16e^{2}}{B}\right),$
where $\rho=\frac{K}{n}$ for subbagging and $\rho=1-(1-\frac{1}{n})^{K}$ for bagging.
This is adapted from and follows immediately from Soloff et al., 2024a (, Theorem 17).
Recall that $\varepsilon$ determines how much we inflate the argmax, so when it’s smaller, the inflated argmax is closer to the standard argmax. Accordingly, when $\varepsilon$ is smaller, Theorem 1 gives a larger $\delta$, reflecting a weaker stability guarantee. Further, note that the instability level $\delta$ decreases (i.e., stability increases) by decreasing the bag size $K$ and increasing the number $B$ of bags.
The size of our model class $|M^{+}|$ has only a small influence on the level of instability of our procedure. For example, when there are two models, $|M^{+}|=2$, the factor $1-\frac{1}{|M^{+}|}=\frac{1}{2}$, giving a constant-factor boost in stability over the case where $M^{+}$ is countably infinite, where $1-\frac{1}{|M^{+}|}$ becomes $1$.
3 Related work
Returning a set of models to acknowledge inherent uncertainty in model selection is an old idea. Early literature on subset selection for linear regression (Spjøtvoll,, 1972) focused on selecting the true model with high probability. More recent works extend the idea of a model confidence set to more general modeling settings (Hansen et al.,, 2011; Zhang et al.,, 2024). These works have a stochastic model for the data and construct a confidence set containing the best-fitting model with high probability. By contrast, our framework places no stochastic assumptions on the data, instead targeting a more modest goal of stability for the selected set.
There is extensive literature on stability in variable selection; see, e.g., Khaire and Dhanalakshmi, (2022) for a recent overview. Work in this area has overwhelmingly focused on “simple” procedures, returning a single subset of variables, and stability is measured by the degree of similarity of the selected subsets of variables. By contrast, Definition 2.2 is not satisfied by simple procedures unless they return the same model most of the time.
Bagging is an important tool for model selection and variable selection (Breiman, 1996b, ). For example, stability selection thresholds the marginal bootstrap inclusion probabilities of each variable (Meinshausen and Bühlmann,, 2010; Shah and Samworth,, 2013). BayesBag (Bühlmann,, 2014) directly averages posterior distributions over different resamplings of the data. Huggins and Miller, (2023) analyze the accuracy and stability properties of the BayesBag approach specifically for Bayesian model selection.
Our work builds on ideas from stable classification. Soloff et al., 2024a introduces a framework for set-valued classification, combining the inflated argmax with bagging to guarantee stability for any base model selection method.
3.1 Alternative selection rules
We discuss the limitations of some alternative approaches to selection $\mathcal{S}$.
The standard argmax.
The default selection rule selects the best model(s) $\hat{M}\in\mathop{\!\arg\!\max}_{m}\hat{w}_{m}$. ($\hat{M}$ may contain more than one model in the case of tied weights.) If the weights contain near-maximizers, the argmax can be sensitive to small perturbations of $\hat{w}$.
Top-$k$.
[…]
Stability selection.
In the context of variable selection, the stability selection method (Meinshausen and Bühlmann,, 2010) selects a single model based on whether each covariate is selected in at least $\tau$ proportion of bags. Concretely, given a vector of weights $\hat{w}$ indexed by $M^{+}=2^{[d]}$, define the marginal variable probabilities $\hat{\Pi}_{j}=\sum_{\begin{subarray}{c}m\subseteq[d]:j\in m\end{subarray}}\hat{w}_{m},$ for $j=1,\ldots,d.$ Stability selection returns
$\text{ip}_{\tau}(\hat{w})$ =$=\left\{\hat{m}_{\tau}\right\},\text{ where }\hat{m}_{\tau}=\Big{\{}j\in[d]:\hat{\Pi}_{j}\geq\tau\Big{\}},$
for some use-specified threshold $\tau$. Similar to the standard argmax (without ties), $\text{ip}_{\tau}(\hat{w})$ only returns a single model, which can be unstable if some marginal probabilities $\hat{\Pi}_{j}$ are near the threshold $\tau$. It also fails to adapt to relationships among variables (§ 4.1).
4 Examples: variable selection in linear models
Consider the problem of variable selection in linear regression (a specific type of model selection); this is a classic problem in bioinformatics (Saeys et al.,, 2007), environmental studies (Effrosynidis and Arampatzis,, 2021), psychology (Vrieze,, 2012), and more.
In this setting, the data is modelled as the form
$y_{i}=\langle x_{i},\beta\rangle+\eta_{i},\ \eta_{i}\sim N(0,\sigma^{2}),$ (6)
where $\sigma^{2}\in\mathbb{R}$ is a vector containing the response error variance, and $\beta\in\mathbb{R}^{d}$ is the vector of regression coefficients. In many practical settings, the true population $\beta$ is hypothesized to be sparse, and the goal is to accurately estimate which elements of $\beta$ are nonzero.
Consider the following common model selection approach: first solve the least absolute shrinkage and selection operator (LASSO, Tibshirani, (1996)) problem
$\hat{\beta}=\text{argmin}_{\beta}\left\{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\langle x_{i},\beta\rangle)^{2}+\lambda R(\beta)\right\},$ (7)
where $\lambda$ is a hyperparameter controlling the sparsity of $\beta$, and the regularization term $R(\beta)=||\beta||_{1}$. Then select variables via
$\hat{m}(\mathcal{D})=\mathbbm{1}\{|\hat{\beta}|>0\}$ (8)
where $\hat{\beta}$ implicitly depends on the dataset $\mathcal{D}$. This is an instance of the setting in (2). In this context, the LASSO provides unstable model selections, particularly in settings with nontrivial correlations among covariates (Meinshausen and Bühlmann,, 2010). Moreover, since many variable selection algorithms, including the LASSO, only return a single set of selected covariates, these methods do not directly provide an approach for assessing uncertainty in the selected model.
4.1 Linear regression with correlated covariates
Consider the setting in which we have a set of covariates $x=\begin{bmatrix}x^{(1)},x^{(2)},\cdots,x^{(d)}\end{bmatrix}^{\top}$ where
$\mathbb{E}[Y|X=x]=x^{(1)}+x^{(3)},$ (9)
and we seek to learn this relationship using sparse variable selection. If $x^{(1)}$ is highly correlated with $x^{(2)}$, and $x^{(3)}$ is highly correlated with $x^{(4)}$ and $x^{(5)}$ (see covariance matrix in Figure 1), then subtle changes in the dataset could result in selecting any of the following models $m_{k}(\mathbf{x})$:
$m_{1}(\mathbf{x})=$$\{x^{(1)},x^{(3)}\},$ $m_{2}(\mathbf{x})=$$\{x^{(1)},x^{(4)}\},$ (10)
$m_{3}(\mathbf{x})=$$\{x^{(1)},x^{(5)}\},$ $m_{4}(\mathbf{x})=$$\{x^{(2)},x^{(3)}\},$
$m_{5}(\mathbf{x})=$$\{x^{(2)},x^{(4)}\},$ $m_{6}(\mathbf{x})=$$\{x^{(2)},x^{(5)}\}.$
If each of the six models above were selected with approximately equal frequency across bags, the output of Algorithm 1, would be approximately
$\hat{w}\approx(1/6)\begin{bmatrix}1,1,1,1,1,1,0,0,\cdots\end{bmatrix}.$ (11)
A selection method that finds the argmax of $\hat{w}$ would be unstable despite the stability of $\hat{w}$ derived from bagging. We seek a method that outputs models corresponding to selected variable sets
$\left\{\{x^{(1)},x^{(2)}\}\times\{x^{(3)},x^{(4)},x^{(5)}\}\right\},$ (12)
accurately indicating that our method suggests either $x_{1}$ or $x_{2}$ should be included in the selected model, but not necessarily both. In other words, this output reflects the fact that our model selection method cannot reliably distinguish among the six models in (10).
Limitations of the argmax.
The argmax applied to $\hat{w}$ in (11) will be highly sensitive to small perturbations in $\hat{w}$. Furthermore, the single best model does not reflect the inherent uncertainty in the model selection and that several models are almost equally likely.
Limitations of top-$k$.
If the user chooses $k=2$, only 2 models will be returned to the user, which does not reflect that any of the 6 models are approximately equally likely. This failure to adapt to uncertainty also applies to the other direction: if the user chooses $k>6$, top-$k$ will still give back $k$ models, despite the fact that only 6 are highly likely.
Limitations of stability selection
The stability selection method (Meinshausen and Bühlmann,, 2010) will select the model corresponding to selected variables $\{x^{(1)},x^{(2)},x^{(3)},x^{(4)},x^{(5)}\}$ if $\tau$ is sufficiently small. This result can be misleading, in that it ultimately selects a model with five selected covariates as opposed to indicating that we lack certainty about which of multiple two-covariate models is correct.
Benefits of the inflated argmax.
For a large enough choice of $\varepsilon$, the $\mathop{\!\arg\!\max}^{\varepsilon}$ would return all 6 models in (10). This output can be reported as a Boolean expression as in (12), which accurately reflects the uncertainty in the model selection.
4.2 Sparse identification of nonlinear dynamical systems (SINDy)
One application of sparse regression, highlighted in our experiments, is the discovery of governing equations of a system from observations. Specifically, we seek to learn the governing equations of a system of the form
$\frac{d}{dt}u(t)=f(u(t)),$ (13)
where $u(t)=\begin{bmatrix}u^{(1)}(t),u^{(2)}(t),\cdots,u^{(d)}(t)\end{bmatrix}^{\top}$ is a $d$-dimensional state vector containing quantities of interest at time $t$, and $f$ is the function describing how the derivative of $u(t)$ evolves with time. The Sparse Identification of Nonlinear Dynamical Systems (SINDy, Brunton et al., (2016)) algorithm represents a differential equation of the form (13) as a weighted sum of terms in a library of functional forms, formulates the problem as an instance of linear regression (6), and solves (7) or a related method to “discover” the differential equations $f$ from data. Specifically, $y_{i}$ in (6) may correspond to estimates of $\frac{d}{dt}u(t_{i})$ at the $i$-th observation time, and $x_{i}$ may correspond to a vector of functionals of $u(t_{i})$, with elements such as $(u^{(1)}(t_{i}))^{3}$ or $u^{(1)}(t_{i})u^{(3)}(t_{i})$.
Bagging Ensemble SINDy (E-SINDy, Fasel et al., (2022)) creates an ensemble of model fits, and with this ensemble, the final model is chosen. This approach corresponds to applying stability selection (Meinshausen and Bühlmann,, 2010; Shah and Samworth,, 2013) to SINDy, using sequentially thresholded ridge regression (STRidge, Rudy et al., (2017)) instead of the LASSO for ${\cal A}$. E-SINDy aggregates the coefficients $\hat{\beta}^{(b)}$ by selecting covariates whose inclusion probability is greater than some prespecified value $\tau\in(0,1)$. (We denote this selection procedure as ipτ in our experiments.) Of the variables with an inclusion probability greater than $\tau$, the $\hat{\beta}$ are set to the average fitted values across bags. In our experiments, we refer to this ensemble approach as $\tilde{\mathcal{A}}_{K,B}$.
5 Experiments
Our experiments assess the stability of five different variable selection approaches, described in Table 1, in two experimental settings.
NOTATION SELECTION CRITERION
$\mathop{\!\arg\!\max}\circ\ \mathcal{A}$ Model(s) with maximum weight assigned by $\mathcal{A}$
$\mathop{\!\arg\!\max}\circ\ \tilde{\mathcal{A}}_{K,B}$ Model(s) with maximum weight assigned by $\tilde{\mathcal{A}}_{K,B}$
top-$k$ $\circ\ \tilde{\mathcal{A}}_{K,B}$ $k$ models with $k$ largest weights assigned by $\tilde{\mathcal{A}}_{K,B}$
$\mathop{\!\arg\!\max}^{\varepsilon}$ $\circ\ \tilde{\mathcal{A}}_{K,B}$ $\varepsilon$-inflated argmax of $\tilde{\mathcal{A}}_{K,B}$
ipτ $\circ$ $\tilde{\mathcal{A}}_{K,B}$ Model with variables with inclusion probability $\geq\tau$ in $\tilde{\mathcal{A}}_{K,B}$
Table 1: Table of notation for each model selection algorithm. $\mathcal{A}$ refers to a base algorithm mapping a dataset $\mathcal{D}$ to empirical predicted probabilities $\hat{w}$ for each possible subset, and $\tilde{\mathcal{A}}_{K,B}$ is the subbagged version of $\mathcal{A}$ for bags with $K$ randomly selected (without replacement) samples.
We perform experiments in two settings: (a) identifying important variables in a synthetic dataset that contains a small collection of highly correlated coefficients, of which only a few are actually used to construct the response $y$ (§ 4.1) and (b) identifying the latent Lotka-Volterra governing differential equations from noisy trajectory data generated from these equations (§ 4.2). For each experimental setting, we generate independent datasets for $N$ trials, each contaminated with independent Gaussian noise, allowing us to examine the stability across a collection of trials. See supplement for details.
In each experiment, we compute sparse estimates of $\beta$ using both a base algorithm $\mathcal{A}$ and a subbagged version of $\mathcal{A}$ denoted $\tilde{\mathcal{A}}_{K,B}$. Based on (5), we compute the stability for trial $j\in[N]$ using
$\delta_{j}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\{\hat{M}_{j}\cap\hat{M}^{\setminus i}_{j}=\varnothing\},$ (14)
This definition of stability measures, for trial $j$, the proportion of LOO selected models $\hat{M}_{j}^{\setminus i}$ for $i\in[n]$ that have no overlap with the set of selected models $\hat{M}_{j}$ returned by training on the entire dataset. The left column of Figure 2 plots the empirical cumulative distribution function (CDF)
$\frac{1}{N}\sum_{j=1}^{N}\mathbbm{1}\{\delta_{j}\leq\delta\},$ (15)
as a function of $\delta\in[0,1]$ to highlight variation across trials with different random seeds. Curves higher on the plot are better in the sense that the corresponding method achieves model selection stability $\delta$ (Definition 2.2) for a larger fraction of trials, and our primary interest is in small $\delta$ (i.e., high stability).
We additionally compute utility-weighted accuracy, inspired by Zaffalon et al., (2012), using
$\frac{1}{N}\sum_{j=1}^{N}\frac{\mathbbm{1}\{m^{*}\in\hat{M}_{j}\}}{|\hat{M}_{j}|},$ (16)
where $m^{*}$ is the “true” data generating model. This measure down-weights the accuracy by the number of models returned $|\hat{M}_{j}|$. This metric is used to compare methods in the center column of Figure 2.
5.1 Penalized regression with LASSO
Figure 1: Covariance matrix for each simulated dataset. Covariate indices 0 and 2 were used to construct the response $y$.
(a) Regression results. In the left plot, we plot the stability curves for the hyperparameters $\tau=0.3$, $k=2,$ and $\varepsilon=0.8$ since these values yield a utility-weighted accuracy of approximately 0.18 for each $\mathcal{S}$.
(b) Lotka-Volterra results. In the left plot, we plot the stability curves for the hyperparameters $\tau=0.63$, $k=2,$ and $\varepsilon=0.09$ since these values yield a utility-weighted accuracy of approximately 0.3 for each $\mathcal{S}$.
Figure 2: (Left column) Empirical CDF for the stability measures $\delta_{j}$, computed in (14), across $j\in[N]$ for each $\mathcal{S}$. (Center column) Utility-weighted accuracy, computed in (16), versus the worst-case instability across $\delta_{j}$. (Right column) Median number of models returned across $j$ versus the maximum $\delta_{j}$ across $j\in[N]$ for each $\mathcal{S}$. (Center and right columns) We plot the values $k\in[1,\dots,6]$, and $\epsilon\in(0,1)$ and $\tau\in(0,1)$ with approximately evenly spread values across their supports.
Our first experiment is based on the motivating example in (9), where we have two important covariates and a small collection of highly correlated covariates. Figure 1 visualizes the covariance structure of the $20$ variables in the design matrix $X$, where the nonzero covariance between distinct covariates is 0.99. For each of the $N=200$ trials, we generate a dataset with $d=20$ covariates and $n=30$ training samples.
Our learning algorithm $\mathcal{A}$ is the LASSO, which corresponds to solving (7) with $R(\beta)=||\beta||_{1}$, and $\tilde{\mathcal{A}}_{K,B}$ corresponds to a subbagged LASSO with $B=10,000$ bags and $K=25$ training samples per bag. We provide additional implementation details about this experiment in the supplement.
Figure 2(a) provides an empirical CDF of $\delta$ (15), the utility-weighted accuracy (16) versus the empirical worst-case instability, and the median number of sets returned across $N$ trials versus the worse-case instability. The empirical CDF across our chosen selection methods $\mathcal{S}$ show that the selection methods that return a single model, the argmax and ipτ, are the least stable. Subagging, which helps in stabilizing the estimated weights $\hat{w}$ assigned to each model $m$, yields marginal gains in stability when using the argmax. Top-$k$ and the inflated argmax, on the other hand, provide enhanced stability due to the fact that these methods are able to return more than one model.
By plotting the utility-weighted accuracy compared to the worst-case instability, we hope to see a method that provides high utility-weighted accuracy while still providing stable model selections. The ipτ selection method fails to control the worst-case instability across a variety values of $\tau$, but is able to achieve good utility-weighted accuracy. Top-$k$ is able to achieve good utility-weighted accuracy with stable estimate. However, the inflated argmax is able to attain similar stability as top-$k$, while achieving a higher utility-weighted accuracy across a range of stability levels. This boost in utility-weighted accuracy is due to the adaptivity of the inflated argmax: this method will output a single model as often as possible, whereas top-$k$ will always output $k$ model selections, leading to a comparatively lower utility in many cases.
We confirm that the inflated argmax frequently outputs a single model by plotting the median number of models returned compared to the worst-case instability across various choices of $\varepsilon$ compared to various choices of $k$.
Figure 3: Top-10 selected models with their corresponding weights $\frac{1}{N}\sum_{j=1}^{N}\hat{w}_{m}^{(j)}$, which averages the weights across trials. Selected covariates for each model are shown with a black square.
To illustrate that returning multiple models, when necessary, can be represented compactly, we borrow notation from logic to express an example of 6 models, reported in Figure 3. Figure 3 shows that the model weights averaged across trials $\left(\frac{1}{N}\sum_{j=1}^{N}\hat{w}^{(j)}\right)$. The top 6 models can be reported as
$\left\{\{x^{(1)},x^{(2)}\}\times\{x^{(3)},x^{(4)},x^{(5)}\}\right\}$
which succinctly describes the model options.
5.2 Lotka-Volterra experiment
Figure 4: (Left) Lotka-Volterra trajectories for one simulated dataset. The time span ranges from 0 to 16, and there are 21 time points with predator-prey information contaminated with $\mathcal{N}(0,0.05)$ distributed noise. (Right) Empirical correlation matrix for the variable library input into SINDy and E-SINDy. We use the library $(c,u^{(1)},(u^{(2)})^{2},u^{(1)}u^{(2)},u^{(2)},(u^{(2)})^{2})$, where $c$ is a constant.
This experiment focuses on discovering Lotka-Volterra dynamics from data using SINDy. The Lotka-Volterra equations were first derived to model chemical reactions in the seminal work (Lotka,, 1910) and later studied in the now classic use case of predator-prey dynamics (Lotka,, 1925). The Lotka-Volterra equations are
$\dot{u}^{(1)}(t)$ $=\;\;\,\alpha u^{(1)}(t)-\beta u^{(1)}(t)u^{(2)}(t),$ (17a)
$\dot{u}^{(2)}(t)$ $=-\gamma u^{(2)}(t)+\zeta u^{(1)}(t)u^{(2)}(t),$ (17b)
where $u^{(1)}$ is the population density of prey, $u^{(2)}$ is the population density of predators, $\dot{u}^{(1)}$ and $\dot{u}^{(2)}$ are the population growth rates, $t$ corresponds to time. The parameters $\alpha,\beta,\gamma,\zeta>0$ respectively refer to the prey’s growth rate, the predator’s impact on the prey’s death rate, the predator’s death rate, and the prey’s impact on the predator’s growth rate (Lotka,, 1925).
We generated synthetic datasets containing $n=21$ time points of abundance data for predator and prey populations, solved from the Lotka-Volterra equations and contaminated with Gaussian noise. $N=100$ trials were independently generated according to the same data generation process. We provide further details of this data generation process in the supplement. We visualize one set of trajectories in Figure 4.
We now imagine we do not know our system is governed by the Lotka-Volterra system and wish to discover the governing model from data using SINDy. To construct our library of functions, we selected degree-two polynomials as our covariates, and for one dataset, we visualize the corresponding covariate correlations in Figure 4. As seen in the correlation matrix, the covariate structure contains highly correlated variables (e.g., $u^{(1)}$ and $(u^{(2)})^{2}$), which impacts the model selection stability of SINDy.
In our experiments, we solve the optimization in (7) with the STRidge penalization (Rudy et al.,, 2017) as our base algorithm $\mathcal{A}$, as in Fasel et al., (2022).
Figure 2(b) shows that (a) selection methods that return a single model (argmax and ipτ, even with subagging the base algorithm) are less stable than methods that return at least one selected model (the inflated argmax and top-$k$), (b) the inflated argmax provides higher utility-weighed accuracy scores for similar worst-case stability levels as top-$k$, and (c) the inflated argmax provides smaller sets of returned models compared to top-$k$ for a wide range of worst-case stability values.
6 Discussion
The combination of bagging and the inflated argmax described in this paper can be applied to any (potentially unstable) black-box model selection procedure and offer stability guarantees described in Theorem 1. We define a notion of stability that is distinct from other works, one that computes the proportion of LOO selected models are disjoint from the selection made by training on the full data. This is in contrast to, for example, covariate-level stability in linear models.
Our experiments illustrate that, in addition to providing theoretical stability guarantees, our method outperforms ad hoc procedures. Specifically, choosing the most frequently selected model across bags provides limited stability gains relative to an unstable base algorithm. Selecting the top-$k$ most frequently selected models across bags enhances stability but is not adaptive to the underlying empirical uncertainty of the model selection process; $k$ models will be returned to the user regardless of the data. Choosing the returned model based on marginal inclusion probabilities of individual covariates is limited to covariate stability tasks and presents hyperparameter selection challenges. In contrast, the inflated argmax adapts to the number of sets returned based on model selection uncertainty, provides an approach for achieving a desired level of stability, and provides competitive utility-weighted accuracies.
Our framework provides multiple opportunities for further investigation, including exploring the interplay of stable model selection and identifiability and opportunities for the stabilization of black-box model selection methods without bagging (which may present computational barriers in some settings).
Acknowledgements
We gratefully acknowledge the support of the National Science Foundation via grant DMS-2023109 and the NSF-Simons National Institute for Theory and Mathematics in Biology via grants NSF DMS-2235451 and Simons Foundation MP-TMPS-00005320. MA gratefully acknowledges the support of the NSF Graduate Research Fellowship Program under Grant No. DGE-1746045. JS gratefully acknowledges the support of the Office of Naval Research via grant N00014-20-1-2337, and the Margot and Tom Pritzker Foundation.
(a) $\varepsilon=0.00$
(b) $\varepsilon=0.15$
(c) $\varepsilon=0.30$
Figure 5: Visualization of the inflated argmax for increasing values of $\varepsilon$. At each corner of the triangle, $m_{1}$, $m_{2}$, and $m_{3}$ are three candidate models to select. When $\varepsilon=0$, this selection process is the same as the standard argmax. If a weight vector $\hat{w}$ lands in one of the colored regions, the inflated argmax returns the associated $\hat{M}$. For example, if $\hat{w}$ lands of the region where red and yellow overlap, models $\{m_{1},m_{2}\}$ are returned.
Appendix A Visualizing the inflated argmax
Figure 5 visualizes the model selection landscape for 3 candidate models for various choices of $\varepsilon$. As $\varepsilon$ increases, the inflated argmax is more likely to return more than one model, evidenced by the larger overlapping areas. Similarly, when $\varepsilon$ decreases, the inflated argmax becomes less and less likely to return more than one set, and in the limit when $\varepsilon\to 0$, the inflated argmax approaches the behavior of the standard argmax.
Appendix B Experiment details
In both our experiments, we utilize a cluster computing system to distribute parallel jobs across CPU nodes.
B.1 Regression example
This section provides further details about the experiment in § 5.1, including the data generation process and a justification for the regularization hyperparameter $\lambda$.
B.1.1 Data generation
In our standard regression example, we simulate datasets for $j\in[N]$ trials, where $N=200$, according to the following process:
$x_{i}\sim\mathcal{N}(0,C),$ (18)
where $x_{i}\in\mathbb{R}^{d}$, $\mathcal{N}$ denotes a Normal distribution, and $C\in\mathbb{R}^{d\times d}$ is visualized in Figure 1. The nonzero off-diagonal elements of $C$ are 0.99, corresponding to extremely correlated covariates. For a single trial $j$, we simulate a dataset with $d=20$ covariates and $i\in[n]$ ($n=30$) samples. We take the first and third columns as the covariates used to create the response $y$,
$y_{i}=x_{i}^{(1)}+x_{i}^{(3)}+\upsilon_{i},\quad\upsilon_{i}\sim\mathcal{N}(0,0.3^{2}),$ (19)
where $\beta=[1,1]^{\top}$, $x_{i}^{(1)}$ and $x_{i}^{(3)}$ are respectively the first and third covariates for sample $i$, and $\upsilon_{i}$ represents independent Gaussian distributed observation noise with mean 0 and standard deviation $0.3$ for sample $i$.
We solve each optimization problem (i.e., fits on the full datasets $\mathcal{D}$, LOO datasets $\mathcal{D}^{\setminus i}$ for $i\in[n]$, and bags $\mathcal{D}^{b}$ for $b\in[B]$ across trials $j\in[N]$) for $\hat{\beta}$ using the optimization function in (7) for $R(\beta)=||\beta||_{1}$, which is equivalent to the LASSO.
B.1.2 Optimization hyperparameter
Selecting $\lambda$ via cross-validation leads to much denser models than the ground truth model, so we increase $\lambda$ to a level that induced more sparsity. This approach implicitly assumes we have prior knowledge of approximately how many coefficients are important in the regression task, but not necessarily which are important. We select $\lambda=0.5$ since this penalization level led to highly sparse variable selections. We note, however, that the theory behind the inflated argmax is model agnostic, meaning that stability guarantees apply for any choice of $\lambda$.
B.2 Lotka-Volterra simulations
This section provides further details of the experiment in § 5.2. We utilize the pysindy Python package (Kaptanoglu et al.,, 2022; de Silva et al.,, 2020) for their implementations of these methods to perform out Lotka-Volterra experiments with SINDy.
B.2.1 Data generation
We generate data from the Lotka-Volterra differential equations described in (17) for $(\alpha,\beta,\gamma,\zeta)=(2/3,4/3,1,1)$ for 21 evenly spread time points $t$ in the range $t\in[0,16]$ and initial condition $[1,1]^{\top}$. Solutions of this ordinary differential equation were obtained with the LSODA (Petzold,, 1983) solver. For each $t\in(0,16]$, non-inclusive of the initial condition, we contaminate each $u^{(1)}(t)$ and $u^{(2)}(t)$ with $\mathcal{N}(0,0.05^{2})$ Gaussian noise, where 0.05 is the standard deviation of the measurement noise level.
To obtain sparse solutions of the regression problem and therefore estimate the governing equations, we solve the optimization problem with STRidge (Rudy et al.,, 2017), which iteratively (a) solves the equation in (7) with $R(\beta)=||\beta||^{2}$, corresponding to solving Ridge regression (Hoerl and Kennard,, 1970), and (b) thresholds $\hat{\beta}$ based on $\omega$, where $\omega$ is the minimum magnitude needed for a coefficient in the weight vector. Any coefficient with an estimated magnitude below $\omega$ is set to 0. This process is repeated until a convergence criterion is met.
B.2.2 Cross-validating optimization hyperparameters
We perform a grid search across two parameters to find a combination that leads to a low validation MSE. The 5-fold validation MSE is measured as
[…]
where $\mathcal{V}(v)$ is the collection of time points included in the validation evaluation for fold $v\in 1,2,\dots,5$; $\hat{u}(t;\lambda,\omega)$ is a vector of solutions, containing $\hat{u}^{(1)}(t;\lambda,\omega)$ and $\hat{u}^{(2)}(t;\lambda,\omega)$, to the model selected by SINDy using optimization parameters $\lambda$ and $\omega$ at time $t$; $u(t)$ is a vector of the observed values for $u^{(1)}(t)$ and $u^{(2)}(t)$ at time $t$. The trajectories were temporally split into 5 disjoint sets, where each set contains temporally adjacent time points.
$\omega$
$\lambda$ 0.16 0.17 0.18 0.19 0.20 0.25
0.0075 0.0538 0.0538 0.0539 0.0982 0.0963 0.1156
0.01 0.0538 0.0538 0.0538 0.0981 0.0963 0.1942
0.02 0.1285 0.1285 0.1287 0.1275 0.1256 0.4403
0.03 0.1285 0.1285 0.3160 0.3166 0.3202 0.3451
0.15 0.0788 - - - -
1 - - - 10.65 - 14.94
Table 2: Table of hyperparameter combinations and corresponding validation MSEs for choosing the STRidge $\lambda$ and $\omega$ hyperparameters. Any combination of hyperparameter values that led to a null model (selecting none of the covariates to include in the model) for any of the validation folds is reported as “-”. We choose the hyperparameter combination that should give sparser models (larger $\lambda$ and larger $\omega$) in the case of tied validation MSE, which leads us to choose $\lambda=0.01$ and $\omega=0.18$ (bolded and highlighted).
Figure 6: Lotka-Volterra solutions (solid lines) for the equations selected using SINDy with hyperparameters $\lambda=0.01$ and $\omega=0.18$ using one observational dataset (points).
We provide a table of validation MSEs averaged across the 5 folds in Table 2. Given our discretization of $\lambda$ and $\omega$ values, we choose $\omega=0.18$ and $\lambda=0.01$ and keep these values fixed in our experiments presented in § 5.1. Figure 6 visualizes the corresponding model selected by 5-fold cross validation. The best estimated model is
$\dot{u}^{(1)}=0.636u^{(1)}+-1.255u^{(1)}u^{(2)},$ (21)
$\dot{u}^{(2)}=-0.839u^{(2)}+0.833u^{(1)}u^{(2)},$ (22)
which selects the correct terms in each differential equation with low error on the fitted $\hat{\beta}$ coefficients.
Appendix C Stability with varying numbers of bags
Figure 7: Empirical CDF of instability measures $\delta_{j}$ across trials $j\in[N]$ in the regression experiment for varying number of bags $B$ across four different selection methods: the $\mathop{\!\arg\!\max}$, top-2, $\mathop{\!\arg\!\max}^{0.8}$, ip0.3. The values of $k$, $\varepsilon$, and $\tau$ were chosen to be consistent with the CDF plot in Figure 2(a).
Figure 8: Empirical CDF of instability measures $\delta_{j}$ across trials $j\in[N]$ in the Lotka-Volterra experiment for varying number of bags $B$ across four different selection methods: the argmax, top-2, argmax0.09, ip0.63. The values of $k$, $\varepsilon$, and $\tau$ were chosen to be consistent with the CDF plot in Figure 2(b).
The computational complexity of estimating the weights associated with each possible model $m\in M^{+}$ is not determined by the cardinality of the set $M^{+}$, but rather the number of bags $B$ used to estimated the weights associated with each model. Intuitively, we can generally expect that a larger number of bags (i.e., larger $B$) will correspond to a better estimate of the weights $\hat{w}$ for each model $m$ across selection methods. Figures 7 and 8 visualize the empirical CDF of instabilities $\delta_{j}$ across trials $j\in[N]$, where $N=200$ in the regression experiments and $N=100$ in the Lotka-Volterra experiment. A separate line is plotted for each choice of $B$.
In both plots, we see that $B=10$ provides the least stability, evidenced by the larger proportion of $\delta_{j}$s near $\delta=1$. The regression example provides evidence that generally, increasing the number of bags improves stability. However, this plot does not appear to clearly evidence this hypothesis as well as the Lotka-Volterra experiment. We hypothesize that since the covariates are highly correlated in the regression example with many reasonable models able to accurately predict the response, this setting may yield more variable stability results across the number of bags compared to the Lotka-Volterra example, which has a clear, unique model.
Appendix D Supplementary regression experiment
To illustrate the utility of the inflated argmax in a setting where this method is particularly useful, we construct an example where the model weights $\hat{w}$ estimated via bootstrapping suggest a high level of uncertainty in the model selection. We focus on one trial, so $N=1$. In constructing this example, we follow a similar procedure as described in (18) and (19). We set $d=200$, $n=300$, $B=10,000$, $K=25$, $\lambda=0.5$, and $\upsilon_{i}\sim N(0,0.5^{2})$, where 0.5 is the response error standard deviation for each sample $i$. The correlation structure of covariates is extended to 200 covariates, where similarly the first 5 covariates are highly correlated with a correlation of 0.99 as in Figure 1, and $x^{(1)}$ and $x^{(3)}$ are used to construct $y$ with $\beta=[1,1]^{\top}$.
Figure 9 shows the selections of among the first 5 covariates out of the 200 total for the top-10 models (covariates 6 through 200 are not selected in any of the top-10 models) using the full dataset versus a LOO dataset. As an illustrative example, we choose a LOO dataset that yielded a different top model compared to the top model using the full dataset.
Figure 9: Models with the top 10 largest weights assigned via subbagging with $B=10,000$ bags for one trial. The y-axis labels correspond to the model weights estimated with 10,000 bags. Covariates within each model (row) that were selected to have a nonzero estimated coefficient by LASSO are shown in black, and covariates that were not selected are shown in white. Only the first 5 covariates are shown since the top 10 models did not select covariates any of the covariates $x^{(6)},x^{(7)},\dots,x^{(200)}$. (Left) Top-10 models using the full dataset where $n=300$. (Right) Top-10 models using a LOO dataset where $n-1=299$.
[…]
The inflated argmax, on the other hand, will select more than one model in this instance for relatively small $\varepsilon$ due to the large uncertainty in the model weights. To concretely show this result, we compute the inflated argmax’s model selection(s) and choose $\varepsilon$ using the result in Theorem 1. Setting $\delta$ to a fixed value, we can compute $\varepsilon$ via the following expression for subbagging
$\varepsilon=\sqrt{\frac{1}{\delta}\left(\frac{1}{n-1}\frac{K/n}{1-K/n}\right)}.$ (23)
Since $|M^{+}|=2^{200}$, we set $\frac{1}{|M^{+}|}$ to 0. Additionally, since the Monte-Carlo error term $\frac{16e^{2}}{B}$ is known to be overly conservative, and since we use a large number of bags $B$ in this experiment, we set this term equal to 0 as well. With $n=300$, $K=25$, and $\delta=0.05$, we compute that $\varepsilon=0.078$.
[…]
$\left\{\{x^{(1)},x^{(2)}\}\times\{x^{(3)},x^{(4)},x^{(5)}\}\right\}.$
This set of returned models is very reasonable given the underlying correlation structure among the covariates.
Moreover, the type of stability in this work only measures changes the stability due to LOO changes to the dataset, which is a relatively small perturbation to the data. In reality, a method would ideally provide robustness to more substantial changes in the dataset (e.g., providing the same, or approximately the same, model selection(s) with a different dataset drawn from the same underlying distribution). The inflated argmax is better equipped than the argmax in this scenario by returning six models, reflecting the large uncertainty in selecting a model.